OpenStudy (anonymous):

Solve the equation to the nearest hundredth of a radian. 2secx-5=0 for 0

OpenStudy (saifoo.khan):

sec x = 1/cosx

OpenStudy (saifoo.khan):

\[\frac{2}{\cos x} = 5 \]Can tou solve now?

OpenStudy (saifoo.khan):

you*

OpenStudy (anonymous):

Umm.. I don't understand how you would find x!

OpenStudy (saifoo.khan):

Do you understand sec x = 1/cosx ?

OpenStudy (anonymous):

Yes, I understand all that!

OpenStudy (anonymous):

2/cosx=5! I understand how u got that too, but now what do I do?

OpenStudy (saifoo.khan):

Cross multiply. you'll get 2/5 = cos x

OpenStudy (anonymous):

how would u find "x" without using a calculator?

OpenStudy (saifoo.khan):

Oh without calc idk, sorry.